# Inertial Magnetic SLAM Systems Using Low-Cost Sensors

### v1

@@ 6.1 Comparison with the MAINS
Long corridor-1 3.53 (0.29) 9.20 (0.91) 0.90 (0.22)
Long corridor-2 2.28 (0.17) 5.23 (1.01) 1.51 (0.16)
Long corridor-3 1.25 (0.13) 2.44 (0.42) 1.08 (0.10)
Corridor-1 5.22 (0.14) 10.46 (0.27) 2.57 (0.14)
Corridor-2 1.93 (0.09) 1.03 (0.77) 1.93 (0.13)
Corridor-3 1.98 (0.03) 1.76 (0.10) 1.78 (0.05)
Spiral staircase-1 3.55 (0.02) 5.22 (0.26) 1.79 (0.00)
Spiral staircase-2 3.59 (0.17) 3.45 (0.12) 2.43 (0.16)
Spiral staircase-3 8.37 (0.20) 9.13 (0.84) 3.20 (0.21)
@@
Table 3: Yaw Error at the end of the Trajectories with a Barometer. Unit: degree
@@
Long corridor-1 1.45 1.59 0.28
Long corridor-2 0.79 1.68 2.34
Long corridor-3 0.29 1.30 1.65
Corridor-1 2.02 2.85 1.77
Corridor-2 3.00 1.38 1.05
Corridor-3 2.19 1.57 0.52
Spiral staircase-1 2.32 1.06 1.83
Spiral staircase-2 0.64 2.35 1.40
Spiral staircase-3 1.58 2.49 0.52
@@
Figure 7: Trajectories (Long corridor-2) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
Figure 8: Trajectories (Corridor-3) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
Figure 9: Trajectories (Spiral staircase-2) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
In the second part of the experiment, the effect of the barometric measurements and inertial sensor quality on the performance of the algorithms was investigated. For the first part, the barometer measurements were removed from all three systems, and the results are summarized in TABLE 4 and TABLE 5. Comparing TABLE 4 with TABLE 2, it can be seen that both the vertical error and horizontal error of all three systems increase slightly without the barometer measurements, but the overall performance remains unchanged. Furthermore, the yaw error of all three systems do not change significantly as seen from TABLE 5. This indicates that the IM-SLAM systems are not heavily reliant on the barometer measurements. However, the barometer measurements are expected to have a more significant impact if explicit loop closure detection is to be performed in the SLAM systems, since the altitude information can help to distinguish different floors in a multi-floor building. For the second part, the IMU measurements quality is varied by averaging the measurements from different numbers of IMU sensors in Osmium MIMU 4844 IMU. The box plots of the horizontal errors at the end of the trajectories vs. the number of IMU sensors are shown in Fig. 10, 11, and 12. It can be seen that the horizontal errors of the tightly-coupled IM-SLAM system are less sensitive to the number of IMU sensors compared to the other two systems, and it outperforms the other two algorithms on the datasets Long corridor and Corridor. On Spiral staircase datasets, the performance of the tightly-coupled IM-SLAM system improves as the number of IMU sensors increases, but it is not as good as MAINSwhen the number of IMU sensors is less than 4. We do not have a clear explanation for this phenomenon yet, but it could be related to the fact that the datasets Spiral staircase have more dynamic motions (e.g., turning and climbing) than the other two types of datasets. Further investigation is needed to fully understand this phenomenon. Lastly, the sudden increase in horizontal error when using 32 IMUsensors is due to exceptionally worse performance on one of the datasets (Spiral staircase-3), which is currently under investigation.
Table 4: Horizontal (Vertical) Error at the End of the Trajectories without a Barometer. Unit: meter
[…]
Long corridor-1 3.80 (1.01) 6.84 (6.75) 0.97 (0.83)
Long corridor-2 2.33 (1.16) 8.41 (1.21) 1.49 (0.11)
Long corridor-3 1.19 (1.21) 4.30 (0.89) 1.23 (0.43)
Corridor-1 5.35 (0.54) 11.05 (0.36) 2.45 (0.28)
Corridor-2 1.82 (0.44) 0.99 (0.95) 1.76 (0.32)
Corridor-3 2.02 (0.52) 1.80 (0.36) 1.89 (0.51)
Spiral staircase-1 3.00 (0.41) 6.90 (1.08) 1.74 (0.16)
Spiral staircase-2 3.39 (0.10) 4.33 (0.90) 2.51 (0.22)
Spiral staircase-3 8.22 (0.65) 8.48 (2.14) 4.13 (0.91)
Table 5: Yaw Error at the End of the Trajectories without a Barometer. Unit: degree
[…]
Long corridor-1 0.87 2.85 0.29
Long corridor-2 1.24 2.17 2.31
Long corridor-3 0.23 1.76 2.00
Corridor-1 2.11 3.12 1.32
Corridor-2 2.90 1.37 0.90
Corridor-3 2.11 1.60 0.28
Spiral staircase-1 2.25 0.23 1.62
Spiral staircase-2 0.56 2.20 0.80
Spiral staircase-3 1.71 0.41 1.40
Figure 10: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Long corridor datasets. The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
@@ 6.2 Ablation Study on the Effect of Barometer and Inertial Sensor Quality
Figure 11: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Corridor datasets. The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
Figure 12: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Spiral staircase datasets. The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
@@ 7 Conclusion
In this paper, a loosely coupled and a tightly coupled IM-SLAMsystem were proposed. Both systems rely solely on commonly available low-cost sensors: an inertial measurement unit (IMU), a magnetometer array, and a barometer. The proposed systems have position errors on the order of meters per 100 meters traveled, which showcases the feasibility of developing magnetic field-based SLAMsystems with low-cost sensors. Experimental results demonstrate that the tightly coupled IM-SLAMsystem achieves better performance than the loosely coupled counterpart. This improvement arises because the tightly coupled system integrates the local and global magnetic field model within a unified framework, rather than employing them in a two-stage manner as in the loosely coupled design. Furthermore, the experiments show that incorporating a barometer primarily benefits IM-SLAMperformance in the vertical direction, and that the IMU quality can be degraded by approximately a factor of 32 without substantially affecting the overall accuracy. These results demonstrate the potential of applying the proposed systems in emergency response scenarios such as mine/fire rescue.

### v2

@@ 6.1 Comparison with MAINS
Long corridor-1 3.53 | 0.29 0.22 | 0.84 0.90 | 0.22
Long corridor-2 2.28 | 0.17 1.52 | 0.60 1.51 | 0.16
Long corridor-3 1.25 | 0.13 0.85 | 0.42 1.08 | 0.10
Corridor-1 5.22 | 0.14 9.39 | 0.27 2.57 | 0.14
Corridor-2 1.93 | 0.09 1.25 | 0.77 1.93 | 0.13
Corridor-3 1.98 | 0.03 1.77 | 0.09 1.78 | 0.05
Spiral staircase-1 3.55 | 0.02 3.46 | 0.19 1.79 | 0.00
Spiral staircase-2 3.59 | 0.17 3.21 | 0.15 2.43 | 0.16
Spiral staircase-3 8.37 | 0.20 10.00 | 0.72 3.20 | 0.21
@@
Figure 2: Trajectories (Long corridor-2) estimated by (a) MAINS, (b) the loosely coupled IM-SLAM system, and (c) the tightly coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
Figure 3: Trajectories (Corridor-3) estimated by (a) MAINS, (b) the loosely coupled IM-SLAM system, and (c) the tightly coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
Figure 4: Trajectories (Spiral staircase-2) estimated by (a) MAINS, (b) the loosely coupled IM-SLAM system, and (c) the tightly coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
In the second part of the experiment, the effect of the barometric measurements and inertial sensor quality on the performance of all systems was investigated. For the first part, the barometer measurements were removed from all three systems, and the results are summarized in TABLE 3. Comparing TABLE 3 with TABLE 2, it can be seen that both the vertical and horizontal errors of all three systems changed slightly (except the loosely coupled IM-SLAM on Long corridor-1) without the barometer measurements, but the overall performance remains unchanged. This indicates that the IM-SLAM systems are not heavily reliant on the barometer measurements. However, the barometer measurements are expected to have a more significant impact if explicit loop closure detection is to be performed in the SLAM systems, since the altitude information can help to distinguish different floors in a multi-floor building. For the second part, the IMU measurement quality is varied by averaging the measurements from different numbers of IMU sensors in Osmium MIMU 4844 IMU. The box plots of the horizontal errors at the end of the trajectories vs. the number of IMU sensors are shown in Fig. 5. It can be seen that the horizontal errors of the tightly coupled IM-SLAM system are less sensitive to the number of IMU sensors compared to the other two systems.
Table 3: Horizontal and Vertical Error at the End of the Trajectories without a Barometer. Unit: meter
@@ 6.2 Ablation Study on the Effect of Barometer and Inertial Sensor Quality
Long corridor-1 3.80 | 1.01 12.80 | 5.23 0.97 | 0.83
Long corridor-2 2.33 | 1.16 1.07 | 0.82 1.49 | 0.11
Long corridor-3 1.19 | 1.21 1.02 | 0.54 1.23 | 0.43
Corridor-1 5.35 | 0.54 10.53 | 0.42 2.45 | 0.28
Corridor-2 1.82 | 0.44 1.20 | 0.95 1.76 | 0.32
Corridor-3 2.02 | 0.52 1.80 | 0.38 1.89 | 0.51
Spiral staircase-1 3.00 | 0.41 3.85 | 0.93 1.74 | 0.16
Spiral staircase-2 3.39 | 0.10 1.77 | 0.49 2.51 | 0.22
Spiral staircase-3 8.22 | 0.65 9.00 | 2.44 4.13 | 0.91
@@
(a) Dataset: Long corridor
(b) Dataset: Corridor
(c) Dataset: Spiral staircase
@@
Figure 5: Box plots of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors. The box plots in each subfigure from left to right correspond to MAINS, the loosely coupled IM-SLAM system, and the tightly coupled IM-SLAM system.
6.3 Ablation Study on the Effect of Pose Aiding Duration on Positioning Performance
In the third part of the experiment, the effect of pose aiding duration on positioning performance was investigated. The pose aiding duration at the start was set from 0 to 30 seconds, and the horizontal error at the end of the estimated trajectory is calculated, as shown in TABLE 4. It can be seen that all systems have reasonably good positioning performance when the pose aiding duration is more than 20 seconds. Therefore, it is recommended to initialize these systems with at least 20 seconds of pose aiding before use.
Table 4: Horizontal Error at the End of the Trajectories with Different Pose Aiding Time. Unit: meter
Dataset Algorithm 0 s 10 s 20 s 30 s
Corridor-3 MAINS 4.66 1.33 2.05 0.95
IM-SLAM (L.) 6.36 1.99 3.65 1.22
IM-SLAM (T.) 3.12 4.12 3.52 1.62
Long corridor-2 MAINS 10.72 7.51 3.56 2.28
IM-SLAM (L.) 9.03 6.66 1.79 0.87
IM-SLAM (T.) 4.39 0.58 0.57 1.51
Spiral staircase-2 MAINS 1.29 0.93 1.45 2.39
IM-SLAM (L.) 1.70 0.96 1.26 1.30
IM-SLAM (T.) 3.74 3.46 5.11 2.39
6.4 Study on the Effect of Model Switching Time
In the fourth part of the experiment, the effect of model switching time is studied. The value of $D$ is varied from 5 to 1500 and $\infty$, corresponding to model switching times of 0.05 s, 15 s, and no switching, respectively. The resulting horizontal error at the end of the estimated trajectory is reported in Table 5. It can be seen that when the switching time is 1 second, the tightly coupled SLAM system has the best overall positioning performance. Given the walking speed (approximately 1 m/s), the distance between two consecutive measurement updates concerning (19) is approximately 1 m, which is roughly on the same order of the length scale of the approximated GP model (0.7 to 1 m). As magnetic fields at locations one length scale apart are less correlated, the tightly coupled SLAM system is very likely to have “recorded” magnetic field measurements (without affecting navigation states estimation) in the exploration phase and used this information to correct drift when revisiting the same locations.
Table 5: Horizontal Error at the End of the Trajectories with Different Model Switching Time. Unit: meter
Dataset 0.05 s 1 s 5 s 15 s $\infty$ s
Long corridor-1 27.70 0.9 5.52 1.28 3.53
Long corridor-2 7.43 1.51 1.16 1.10 2.28
Long corridor-3 20.70 1.08 1.30 1.10 1.25
Corridor-1 1.87 2.57 2.60 5.83 5.22
Corridor-2 10.11 1.93 2.71 1.89 1.93
Corridor-3 18.61 1.78 0.69 3.06 1.98
Spiral staircase-1 5.48 1.79 1.41 3.55 3.55
Spiral staircase-2 9.42 2.43 2.91 2.53 3.59
Spiral staircase-3 8.02 3.20 3.73 7.99 8.37
6.5 Introducing Pseudo Measurements
In the last part of the experiments, we investigate whether a pseudo measurement can be used to smooth the trajectories in the tightly coupled SLAM system. The measurement equation (19) is augmented with a pseudo measurement,
$y^{\text{\tiny{pseudo}}}_{k}$ $=h^{\text{\tiny{pseudo}}}(x_{k})+e_{k}^{\text{\tiny{pseudo}}}$ (22a)
$h^{\text{\tiny{pseudo}}}(x_{k})$ $=R^{\top}_{k}\nabla\Psi(p_{k})\eta_{k}-\Phi(0)\theta_{k}.$ (22b)
The pseudo measurement is the difference between magnetic fields at the center of the magnetometer array predicted the Gaussian process model and the polynomial model. Ideally, the pseudo measurement should be zero, and the covariance of $e_{k}^{\text{\tiny{pseudo}}}$ reflects the strictness of this constraint.
This pseudo measurement was introduced in the experiments on the datasets Corridor-3 and Spiral staircase-2. The dataset Long corridor was not selected for this experiment because the trajectory in Figure 2 was less jerky. The covariance of $e_{k}^{\text{\tiny{pseudo}}}$ was set to $0.005I_{3}$ and $0.01I_{3}$, respectively. Here, $I_{3}$ denotes the identity matrix. The trajectories reported by the tightly coupled IM-SLAM system after introducing the pseudo measurement are shown in Figure 6. Comparing the trajectories in Figure 3-(c) and Figure 4-(c), it can be seen that the trajectories in Figure 6 are smoother.
Figure 6: These two figures (a) Trajectory (Corridor-3) (b) Trajectory (Spiral Staircase-2) illustrate the smoothing effect after introducing the pseudo measurement in the tightly coupled IM-SLAM system.
@@ 7 Conclusion
In this paper, loosely coupled and tightly coupled IM-SLAM systems are proposed. Both systems rely solely on commonly available low-cost sensors: an inertial measurement unit, 30 magnetometers, and a barometer. The proposed systems have position errors on the order of meters per 100 meters traveled, which showcases the feasibility of developing magnetic-field-based SLAM systems with low-cost sensors. Experimental results demonstrate that the tightly coupled IM-SLAM system achieves better performance than the loosely coupled counterpart. Furthermore, the experiments show that incorporating a barometer primarily benefits IM-SLAM performance in the vertical direction, and that the IMU quality can be degraded by approximately a factor of 32 without substantially affecting the overall accuracy of the tightly coupled SLAM system. These results demonstrate the potential of applying the proposed systems in emergency response scenarios such as mine or fire rescue, where one cannot rely on a GNSS or visual-based localization system.